MODIS Today: Status

Warning: file_get_contents(/webdata/web/google-earth/ge/htdocs/modis-today/templates/../status-messages/banner.html): failed to open stream: No such file or directory in /webdata/web/google-earth/ge/htdocs/modis-today/templates/status-messages.tpl.php on line 18

Aug 10, 2016 The Aqua anomaly affecting the time stamp of MODIS and CERES data has been resolved as of 222/20:56:05 UTC. The MODIS and CERES packet time stamps have been verified and are now showing the correct time stamp. Analysis indicates that the time stamps were incorrect for MODIS/CERES from 221/04:49:10 - 222/20:56:05 UTC. At the start of the anomaly, time generation reverted to an epoch time of 1958/001/00:00:00 (YEAR/DOY/HH:mm:ss) and continued to count until the anomaly was resolved.

Oct 12, 2022 Terra MODIS Science data is currently unavailable due to a constellation exit maneuver scheduled to start around October 12, 2022. Please see here for more information.

Feb 14, 2023 Terra MODIS Science data is currently unavailable due to an X-band anomaly that occurred on February 12, 2023.

Aug 9, 2016 The Aqua spacecraft experienced an anomaly that affects the time stamp on MODIS and CERES data. This anomaly started at 221/04:49:10 UTC and has yet to be resolved. The Aqua C&DH Engineer reports: "On DOY 221 Aqua experienced a TIE Reboot, which caused the time for one of the clocks on board the spacecraft to reset to DOY 001, 1958. Because of this all the times for the MODIS and CERES packets are incorrect since the anomaly." EDOS is analyzing the issue.

Mar 3, 2016 We have received information that thermal evissive bands (TEB) have stabilized as of Feb 29. However, for official information, we recommend users monitor the MODIS Characterization Support Team (MCST) web page for the latest information:

Mar 1, 2016 Terra spacecraft entered safe mode at 14:33:17 GMT on 2016/049 during an inclination adjustment maneuver (IAM). This caused the MODIS to enter safe mode as well with the nadir and the space-view doors closed. The spacecraft was recovered from the Safe-Mode to normal mode on 2/22/2016 and subsequently the MODIS instrument was successfully transitioned to science mode on 2/24/2016. No instrument data was available for the period 2016/049 - 2016/054.

Feb 26, 2016 Terra direct broadcast has resumed, but users may notice lingering issues with Infrared channels as the instrument temperatures stabilize.

Feb 25, 2016 Terra direct broadcast has resumed.

Feb 24, 2016 Per Goddard Space Flight Center (GSFC): The MODIS Instrument was configured for science today, February 24, 2016 at approximately 20:30 UTC. The instrument temperatures are stabilizing, but data should again be available via direct broadcast.

Feb 23, 2016 Per Goddard Space Flight Center (GSFC): "Terra spacecraft remained in safe hold throughout the weekend allowing time for the Flight Operations Team to determine cause of entry into safe hold, develop recovery procedures, test against simulator and present to Anomaly Review Board plans for recovery. February 22, 2016, the recovery procedures were approved. The Flight Operation Team was successful in restoring operations to the Spacecraft Computer. Instrument recovery activities will start February 23, 2016 and likely take multiple days. We will send out an update when MODIS Direct Broadcast is available."

Feb 18, 2016 Per Goddard Space Flight Center (GSFC) regarding Terra: "On February 18, 2016 at approximately 1433z, the Terra spacecraft entered safe mode. Spacecraft is currently in a stable configuration however there is no science data being collected. Recovery is likely to take multiple days. Will post message when Terra MODIS is again available via direct broadcast."

Jan 19, 2009 Our X-band antenna is back online and operating normally. We believe some moisture entered a component on the antenna pedestal and froze. The problem resolved itself when the temperature went above freezing last week. Normal data processing has resumed. (LiamG)

Dec 23, 2009 Our antenna is having problems and we are working with SeaSpace to fix it. In the interim we will continue to ingest DB MODIS data from our colleagues at GSFC, USF, RSAC, UT-Austin, and UNM. (LiamG)

Aug 26, 2009 The Terra satellite went into safe mode at 13:56:25 today due to an anomaly with the Science Formatting Equipment. NASA is investigating and will provide more details as they become available. (LiamG)

Jun 24, 2009 Our building experienced a complete power outage this morning for more than 60 minutes. We are recovering our systems slowly. MODIS Today should be up to date in the next 2-3 hours (LiamG)

Mar 11, 2009 After a processing system failure early on Saturday morning, we are now back online with routine data processing from Terra and Aqua. For those who use our KML files in Google Earth, the transition to our new web server is now complete. At the moment, all systems are operating normally. (LiamG)

Mar 03, 2009 Our X-band antenna is offline at the moment, and we are working on a fix. We continue to ingest Level 0 data from other DB sites in the USA, however this backup processing is also experiencing some unrelated problems at this time. We are working to fix this system as well. Finally, we are aware that there are some lingering issues in the Google Earth KML files related to the switchover to our new web server. Please bear with us as these are fixed.

Feb 25, 2009 We have switched to a new Web server for the MODIS Today imagery and associated Google Earth content. Please update your bookmarks. The new server offers 13TB of storage space for both MODIS and AVHRR imagery, and much faster response times than our old server.

Aug 31, 2008 We've had some problems with the MODIS Today images due to changes in the software which does the image processing. We've rolled back to an older version of the software until we can fix the problems. Sorry for the inconvenience.

Jun 13, 2008 Terra direct broadcast has resumed.

Jun 1, 2008 The direct broadcast transmitter on the Terra spacecraft has been turned off while the Flight Operations Team resolves a problem with abnormal transmitter voltages and temperatures. We will continue to process Terra data as it becomes available from the NOAA bent pipe.

Apr 16, 2008 The Terra SFE was reset successfully, and normal DB operations have resumed.

Apr 16, 2008 The TERRA Science Formatting Equipment (SFE) suffered an anomaly today, sometime between 15:00 and 16:00 UTC. The Flight Operations Team is currently evaluating the data.

Apr 14, 2008 Our loaner receiver is installed, but is only receiving Aqua data reliably at present. Terra data are being ingested from other US DB sites in the interim. We are working with SeaSpace to come up with a fix.

Apr 07, 2008 A loaner receiver from SeaSpace Corp. should be installed this morning. If all goes well, we should be back in action later today.

Apr 04, 2008 Our X-band receiver is down at the moment. We expect to have a backup receiver installed on Monday. Until then, we will ingest data from GSFC and OSU to retain our product generation capability. Product latency will be slightly increased as a result.

Dec 13, 2007 The Terra Science Formatting Equipment (SFE) experienced another anomaly this morning at 347/13:11:17 UT. The Flight Operations Team (FOT) is presently investigating the anomaly. The spacecraft and instruments are safe. There is no science data recording or Direct Broadcast at this time. We will keep you posted.

Dec 13, 2007 The Terra operations team executed recovery procedures to return the Terra Science Formatting Equipment (SFE) to nominal operation during a TDRSS contact starting at 20:22 UTC today. The SFE began operating normally at 20:30 UTC. Direct broadcast and science data recording are back to normal.